OpenStudy (anonymous):

An equation is shown below. -2(4x - 1) - 7 = 5 Which statement shows a correct next step in solving the equation? The equation can become -2(4x - 1) = -2 by applying the distributive property. The equation can become -2(4x - 1) = 12 by applying the addition property of equality. The equation can become -2(4x - 1) = 12 by applying the commutative property of addition The equation can become -2(4x - 1) = -2 by applying the subtraction property of equality.

OpenStudy (ranga):

To solve the equation, the first step is to add 7 to both sides: -2(4x - 1) = 5 + 7 = 12 This is the addition property of equality.